OpenStudy (anonymous):

A=(b1=b2)h ----------- 2 for b2 can some1 show d work

OpenStudy (anonymous):

2A = h(b1 + b2) 2A/h = b1 + b2 b2 = (2A/h) - b1

OpenStudy (anonymous):

2A/h+b1=b2

OpenStudy (anonymous):

tnks